OpenStudy (abbycross167):

Can someone please help me with this question? I have no clue how to do it.... Graph the inequality 102c + 150m + 41 > 500

OpenStudy (abbycross167):

@jdoe0001 and @mathstudent55 could one of y'all please try to show me how to graph?

OpenStudy (abbycross167):

@jim_thompson5910 could you try to help me?

jimthompson5910 (jim_thompson5910):

Is this the problem we did a few days ago? If so, then treat 'c' as 'x'. Also, treat 'm' as 'y' so you'll have 102x + 150y + 41 > 500

jimthompson5910 (jim_thompson5910):

to graph `102x + 150y + 41 > 500` you'll first need to graph `102x + 150y + 41 = 500` to form the boundary line

OpenStudy (abbycross167):

No sir/ma'am I don't think it is,

OpenStudy (abbycross167):

@jim_thompson5910 I don't know how to find the x and y values that I need to graph the inequality...

jimthompson5910 (jim_thompson5910):

what is the value of y when x = 0 ?

OpenStudy (abbycross167):

5? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

102x + 150y + 41 = 500 102*0 + 150y + 41 = 500 ... replace x with 0 150y + 41 = 500 now solve for y

OpenStudy (abbycross167):

would I add 150 and 41?

jimthompson5910 (jim_thompson5910):

no, they are on opposite sides of the equal sign

jimthompson5910 (jim_thompson5910):

you need to move the 41 over, so subtract 41 from both sides

OpenStudy (abbycross167):

oh ok

OpenStudy (abbycross167):

so before i subtract it would look like this... 150y + 41-41 = 500-41? or this 41-150y + 41 = 500-41

jimthompson5910 (jim_thompson5910):

`150y + 41-41 = 500-41` is the correct way

OpenStudy (abbycross167):

ok, 150y=459 y=150?

jimthompson5910 (jim_thompson5910):

to isolate y in `150y=459`, you need to divide both sides by 150

jimthompson5910 (jim_thompson5910):

150y means 150 times y division is used to undo the multiplication

OpenStudy (abbycross167):

so it would equal 3? y=3?

jimthompson5910 (jim_thompson5910):

you should have 459/150 = 3.06

OpenStudy (abbycross167):

oh yes sir/ma'am

jimthompson5910 (jim_thompson5910):

ok so if x = 0, then y = 3.06 this is one point on the line you need one more point to graph the line

jimthompson5910 (jim_thompson5910):

plug in x = 1 and solve for y

OpenStudy (abbycross167):

I thought that's what i just did?

jimthompson5910 (jim_thompson5910):

no you plugged in x = 0

OpenStudy (abbycross167):

oh

OpenStudy (abbycross167):

so in this -- 102x + 150y + 41 = 500 I would turn 150y into 3.06? like this... 102x + 3.06+ 41 = 500

jimthompson5910 (jim_thompson5910):

102x + 150y + 41 = 500 102*1 + 150y + 41 = 500 ... replace x with 1 102 + 150y + 41 = 500 now isolate y

OpenStudy (abbycross167):

would I subtract 41 to both sides?

jimthompson5910 (jim_thompson5910):

why not combine like terms on the left side?

OpenStudy (abbycross167):

what do you mean?

OpenStudy (abbycross167):

like-- 150y + 41 = 500+102?

jimthompson5910 (jim_thompson5910):

`102 + 150y + 41 = 500` 102 and 41 are like terms on the left side you can combine them

OpenStudy (abbycross167):

so 150y+143=500?

jimthompson5910 (jim_thompson5910):

yep, now you subtract 143 from both sides to undo the addition

OpenStudy (abbycross167):

so, 143-150y+143=500-143? 7y=357?

jimthompson5910 (jim_thompson5910):

you do NOT say 150y-143 = 7y 150y and 143 are NOT like terms, we cannot combine them

jimthompson5910 (jim_thompson5910):

leave 150y alone as is

OpenStudy (abbycross167):

ok, well then I'm not sure how to do it

jimthompson5910 (jim_thompson5910):

your next step is that you have 150y = 357 now divide both sides by 150

OpenStudy (abbycross167):

357/150=2.38

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so before we found (0,3.06) as one point another point is (1,2.38)

jimthompson5910 (jim_thompson5910):

plot the two points (0,3.06) and (1,2.38) then draw a straight line through them

OpenStudy (abbycross167):

oh ok... thats it?

jimthompson5910 (jim_thompson5910):

that will take care of the line `102x + 150y + 41 = 500` let me know when you've graphed it there's a few more steps after this

OpenStudy (abbycross167):

ok I'm done

jimthompson5910 (jim_thompson5910):

now we must use a test point (0,0) is the easiest since 0 is easy to work with

jimthompson5910 (jim_thompson5910):

(0,0) means x = 0 and y = 0 let's plug them both in 102x + 150y + 41 > 500 102*0 + 150y + 41 > 500 ... replaced x with 0 102*0 + 150*0 + 41 > 500 ... replaced y with 0 0 + 0 + 41 > 500 41 > 500

jimthompson5910 (jim_thompson5910):

now ask yourself this: is `41 > 500` a true inequality? or is it false?

OpenStudy (abbycross167):

false

jimthompson5910 (jim_thompson5910):

since `41 > 500` is false, this means `102x + 150y + 41 > 500` is false when (x,y) = (0,0) so you do NOT shade the region where (0,0) is located. You shade the region on the opposite side of the line

jimthompson5910 (jim_thompson5910):

In this case, it means you shade above the line `102x + 150y + 41 = 500`

OpenStudy (abbycross167):

 So this is how it would look?